OpenStudy (anonymous):

40=-2(m+4)+m WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (austinl):

Solve for m?

OpenStudy (anonymous):

Answer choices: A.-24 B.-16 C.-32 D.

OpenStudy (anonymous):

d.-48

OpenStudy (anonymous):

40 = -2(m + 4) + m 40 = -2m - 8 + m 32 = -m m = -32

OpenStudy (austinl):

You are supposed to guide them to the answer basilimus, not give it to them.